# Thick isotopy property and the mapping class groups of Heegaard splittings

### v4

@@ 4. The min-max theorem
Let $k\in\{1,\ldots,n\}$. Note that by the isotopy extension theorem, we can fix diffeomorphisms $f_{s}:\Sigma\to\Sigma^{k}_{s}$ for $s\in\mathrm{int}\,B$ simultaneously. In the next lemma, which follows from the definition of $\{\Sigma^{k}_{s}\}_{s\in B}$, we identify $\Gamma(\Sigma)$ with $\Gamma(\Sigma^{k}_{s})$ through these diffeomorphisms.
@@
The proof is by contradiction. Suppose that there exists a line segment $\ell\subset B$ connecting 00 with $s\in B\setminus U_{0}$ such that $\ell$ contains a point $u\in U_{0}$. We can find a compressing disk $D$ for $\Sigma^{k}_{u}$ with $\mathrm{diam}\,\partial D\leq\delta$. Let $Q$ be a $3$-ball of diameter $\leq\delta$ that contains $\partial D$. After perturbing $Q$, $\partial Q$ intersects $\Sigma^{k}_{0}=\Sigma$, $\Sigma^{k}_{s}$ and $\Sigma^{k}_{u}$ transversely. If $Q\cap\Sigma_{s}$ contains circles that are essential in $\Sigma^{k}_{s}$, one of such circles bounds a compressing for $\Sigma^{k}_{s}$. As $\Sigma^{k}_{s}$ is $\delta$-locally incompressible, this case cannot occur. Thus, all the circles in $Q\cap\Sigma^{k}_{s}$ are inessential in $\Sigma^{k}_{s}$. Similarly, all the circles in $Q\cap\Sigma$ are inessential in $\Sigma$. By the innermost disk argument, we can isotope $Q$ so that $Q$ is contained in the region between $\Sigma$ and $\Sigma^{k}_{s}$, which is diffeomorphic to $\Sigma\times I$. Note that $Q$ still intersects $\Sigma^{k}_{u}$ so that one of the circles in $Q\cap\Sigma^{k}_{u}$ is essential in $\Sigma^{k}_{u}$. Thus, $\Sigma^{k}_{u}$ is compressible in the product region. This is a contradiction because $\Sigma^{k}_{u}$ is isotopic to a level surface in the product region. ∎
@@
    $B\setminus U_{2}$ is a star-shaped region, where $U_{2}$ is the set of those points $s$ such that $\Sigma_{2s}$ is $\delta$-compressble.
@@
First, note that there is a natural action of $\mathrm{Diff}(M,\Sigma)$ on $\Gamma(\Sigma)$, which iduces the action on $[S^{d-1},\Gamma(\Sigma)]$. By construction, $\{\Sigma_{1s}\}_{s\in B}$ is the image of $\{\Sigma_{0s}\}_{s\in B}$ by $\tilde{h}_{1}$. In other words, $\{\Sigma_{1s}\}_{s\in B}$ can be recovered from $\tilde{h}_{1}\cdot\varphi$ as follows. We repeat the same construction as in Step 1. Take a triangulation $\mathcal{K}$ of $S^{d-1}$ such that $\tilde{h}_{1}\cdot\varphi:S^{d-1}\to\Gamma(\Sigma)$ is simplicial. For each vertex $s$ of $\mathcal{K}$, $\Sigma_{1s}$ is the result of compressing $\Sigma$ across the disk representing $\tilde{h}_{1}\cdot\varphi(s)$. We regard $B$ as a cone over $\mathcal{K}$ and extend the construction of $\Sigma_{1s}$ linearly over each simplex, recovering the sweep-out $\{\Sigma_{1s}\}_{s\in B}$.
@@
By assumption, there is a homotopy $\Phi:[1,2]\times S^{d-1}\to\Gamma(\Sigma)$ such that $\Phi_{1}=\tilde{h}_{1}\cdot\varphi$ and $\Phi_{2}$ is a product of $\varphi_{k}$’s. After passing to a subdivision of $\mathcal{K}$ if necessary, we can extend $\mathcal{K}$ to a triangulation $\mathcal{L}$ of $[1,2]\times S^{d-1}$ ($=[1,2]\times\partial B$) such that $\Phi$ is simplicial with respect to $\mathcal{L}$. Define $\Sigma_{t0}=\Sigma$ for $t\in[1,2]$. For a vertex $(t,s)$ of $\mathcal{L}$, define $\Sigma_{ts}$ as the result of compressing $\Sigma$ across the disk representing $\Phi(t,s)$. Again regard $[1,2]\times B$ as a cone over $\partial([1,2]\times B)$ and extend the construction of $\Sigma_{ts}$ linearly over each simplex, obtaining the sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[1,2]\times B}$.
@@
This sweep-out satisfies the desired property. Indeed, if $\ell\subset B$ is a radius of $B$, $\{\Sigma_{2s}\}_{s\in\ell}$ appears in some $\{\Sigma^{k}_{s}\}_{s\in B}$ as a subfamily. Thus (2) and (3) hold. (1) is obvious from the construction. ∎
@@
After perturbing $Q$, we assume that $\partial Q$ intersects $\Sigma_{y}$ transversely. We can find a circle in $\partial Q\cap\Sigma_{y}$ that is essential in $\Sigma_{y}$. Indeed, if all the circles in $\partial Q\cap\Sigma_{y}$ were inessential, by the innermost disk argument, $\Sigma_{y}$ could be isotoped so that afterward $\Sigma_{y}\subset Q$. This is impossible because $\Sigma_{y}$ is a Heegaard surface and $M$ is not a $3$-sphere. So one of the circles in $\partial Q\cap\Sigma_{y}$ bounds a compressing disk for $\Sigma_{y}$. Define $D_{y}$ as such a disk. By definition, $D_{y}\cap D_{y_{j}}=\emptyset$ for $21\leq j\leq 2^{i+1}-$ 2 and $\mathrm{diam}\,\partial D_{y}<3^{1+\cdots+(i+1)}\cdot\delta$, which prove the claim. ∎
@@
If $d=2$, we can deduce a contradiction as follows. Let $V$ and $W$ be the handlebodies in $M$ bounded by $\Sigma$. Denote by $\Gamma_{V}(\Sigma)$ (resp. $\Gamma_{W}(\Sigma)$) the subcomplex of $\Gamma(\Sigma)$ spaned by compressing disks for $\Sigma$ that lie in $V$ (resp. $W$). Furthermore, denote by $\Gamma_{VW}(\Sigma)$ the union of all simplices that contain vertices in both $\Gamma_{V}(\Sigma)$ and $\Gamma_{W}(\Sigma)$. Thus, $\Gamma(\Sigma)=\Gamma_{V}(\Sigma)\cup\Gamma_{VW}(\Sigma)\cup\Gamma_{W}(\Sigma)$. Recall that the choice of $\varphi$ is arbitrary as long as it is homotopically nontrivial. By Claim 2.7 in [3], we can assume that $\varphi$ is represented by a loop $\gamma$ in $\Gamma(\Sigma)$ with the following properties:

### v5

@@ 4. The min-max theorem
Let $k\in\{1,\ldots,n\}$. Note that by the isotopy extension theorem, we can fix a diffeomorphism $f_{s}:\Sigma\to\Sigma^{k}_{s}$ for $s\in\mathrm{int}\,B$ simultaneously. In the next lemma, which follows from the definition of $\{\Sigma^{k}_{s}\}_{s\in B}$, we identify $\Gamma(\Sigma)$ with $\Gamma(\Sigma^{k}_{s})$ through these diffeomorphisms.
@@
The proof is by contradiction. Suppose that there exists a line segment $\ell\subset B$ connecting 00 with $s\in B\setminus U_{0}$ such that $\ell$ contains a point $u\in U_{0}$. We can find a compressing disk $D$ for $\Sigma^{k}_{u}$ with $\mathrm{diam}\,\partial D\leq\delta$. Let $Q$ be a $3$-ball of diameter $\leq\delta$ that contains $\partial D$. After perturbing $Q$, $\partial Q$ intersects $\Sigma^{k}_{0}=\Sigma$, $\Sigma^{k}_{s}$ and $\Sigma^{k}_{u}$ transversely. If $Q\cap\Sigma_{s}$ contains circles that are essential in $\Sigma^{k}_{s}$, one of such circles bounds a compressing disk for $\Sigma^{k}_{s}$. As $\Sigma^{k}_{s}$ is $\delta$-locally incompressible, this case cannot occur. Thus, all the circles in $Q\cap\Sigma^{k}_{s}$ are inessential in $\Sigma^{k}_{s}$. Similarly, all the circles in $Q\cap\Sigma$ are inessential in $\Sigma$. By the innermost disk argument, we can isotope $Q$ so that $Q$ is contained in the region between $\Sigma$ and $\Sigma^{k}_{s}$, which is diffeomorphic to $\Sigma\times I$. Note that $Q$ still intersects $\Sigma^{k}_{u}$ so that one of the circles in $Q\cap\Sigma^{k}_{u}$ is essential in $\Sigma^{k}_{u}$. Thus, $\Sigma^{k}_{u}$ is compressible in the product region. This is a contradiction because $\Sigma^{k}_{u}$ is isotopic to a level surface in the product region. ∎
@@
    $B\setminus U_{2}$ is a star-shaped region, where $U_{2}$ is the set of those points $s$ such that $\Sigma_{2s}$ is $\delta$-compressible.
@@
First, note that there is a natural action of $\mathrm{Diff}(M,\Sigma)$ on $\Gamma(\Sigma)$, which induces the action on $[S^{d-1},\Gamma(\Sigma)]$. By construction, $\{\Sigma_{1s}\}_{s\in B}$ is the image of $\{\Sigma_{0s}\}_{s\in B}$ by $\tilde{h}_{1}$. In other words, $\{\Sigma_{1s}\}_{s\in B}$ can be recovered from $\tilde{h}_{1}\cdot\varphi$ as follows. We repeat the same construction as in Step 1. Take a triangulation $\mathcal{K}$ of $S^{d-1}$ such that $\varphi:S^{d-1}\to\Gamma(\Sigma)$ is simplicial. For each vertex $s$ of $\mathcal{K}$, choose a compressing disk $D_{s}$ for $\Sigma$ that represents $\varphi(s)$. Then, $\tilde{h}_{1}(D_{s})$ is a compressing disk that represents $\tilde{h}_{1}\cdot\varphi(s)$. Fix an identification $\kappa_{s}:B^{2}_{1}\times[-1,1]\to\tilde{h}_{1}(N(D_{s}))$ for each $s\in\mathcal{K}^{0}$, and define the homotopy $\theta^{s}_{\tau}:M\to M$ to be
$\theta^{s}_{\tau}:=\kappa_{s}\circ\rho_{\tau}\circ\kappa_{s}^{-1}$
on $\tilde{h}_{1}(N(D_{s}))$ and to be the identity on the complement of a small neighborhood of $\tilde{h}_{1}(N(D_{s}))$. On each simplex $\tilde{\sigma}$ of $\widetilde{\mathcal{K}}$ with vertices $0=s_{0},s_{1},\ldots,s_{d}$, define $\Sigma_{1s}$ by the image of the homotopy
$\theta^{s_{1}}_{\tau_{1}}\circ\cdots\circ\theta^{s_{d}}_{\tau_{d}}:M\to M,$
where $s=\tau_{1}s_{1}+\cdots+\tau_{d}s_{d}$ and $\tau_{1},\ldots,\tau_{d}\in I$. Note that the above construction contains an ambiguity regarding the choice of an identification $B^{2}_{1}\times[-1,1]\to\tilde{h}_{1}(N(D_{s}))$: we have to see that $\kappa_{s}$ coincide with the image of the “standard” one, that is $\tilde{h}_{1}\circ\iota_{s}$. But these two maps are isotopic by the uniqueness of tubular neighborhoods, and hence the corresponding sweep-outs can be interpolated via this isotopy. In this way, we recover the sweep-out $\{\Sigma_{1s}\}_{s\in B}$.
@@
By assumption, there is a homotopy $\Phi:[1,2]\times S^{d-1}\to\Gamma(\Sigma)$ such that $\Phi_{1}=\tilde{h}_{1}\cdot\varphi$ and $\Phi_{2}$ is a product of $\varphi_{k}$’s. By the relative simplicial approximation theorem (see e.g. [33]), we can extend $\{1\}\times\mathcal{K}$ to a triangulation $\mathcal{L}$ of $[1,2]\times S^{d-1}$ ($=[1,2]\times\partial B$) such that $\Phi$ is simplicial with respect to $\mathcal{L}$. In the following, we will construct a sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[1,2]\times B}$ that “shadows” the homotopy $\Phi$.
The construction is similar to that in Step 1. Extend the triangulation $\mathcal{L}$ over $\partial([1,2]\times B)$ by adding the two points $(1,0)$ and $(2,0)$ as vertices. Regarding $[1,2]\times B$ as a cone over $\partial([1,2]\times B)$, we obtain a triangulation $\widetilde{\mathcal{L}}$ of $[1,2]\times B$. For each vertex $v=(t,s)$ on $\mathcal{L}$, choose a representative $D_{v}$ of $\Phi(v)$ so that if $v$ and $v^{\prime}$ are in the same simplex, $D_{v}$ and $D_{v^{\prime}}$ are disjoint. For $v=(1,0)$ or $(2,0)$, let $D_{v}=\emptyset$.
As before, fix an identification $\kappa_{v}:B^{2}_{1}\times[-1,1]\to N(D_{v})$ for each vertex $v$ of $\mathcal{L}$, and define $\theta^{v}_{\tau}:M\to M$ to be $\kappa_{v}\circ\rho_{\tau}\circ\kappa^{-1}_{v}$ on $N(D_{v})$ and to be the identity outside a small neighborhood of $N(D_{v})$. It suffices to construct the sweep-out on every simplex $\tilde{\sigma}$ of $\widetilde{\mathcal{L}}$. Let $v_{0},\ldots,v_{d+1}$ be the vertices of $\tilde{\sigma}$ and regard $\tilde{\sigma}$ as a $(d+1)$-cube in Euclidean space with $v_{0}=0$. For $x=\tau_{1}v_{1}+\cdots+\tau_{d+1}v_{d+1}\in\tilde{\sigma}$, $\tau_{1},\ldots,\tau_{d+1}\in I$, define $\Sigma_{x}$ as the image of $\Sigma$ by the composition
$\theta^{v_{1}}_{\tau_{1}}\circ\cdots\circ\theta^{v_{d+1}}_{\tau_{d+1}}.$
Note that for $v\in\{2\}\times S^{d-1}$, the disk $D_{v}$ representing $v$ may not coincide with the standard one which has been chosen in Step 1. Of course, $D_{v}$ is isotopic to the standard choice, and in addition if $v_{1},\ldots,v_{l}$ be vertices adjacent to $v$ and all the $D_{v_{i}}$ have already been in the standard position, then we can isotope $D_{v}$ to the standard position in the complement of $D_{v_{i}}$. Thus, after these isotopies, we may assume that all the $D_{v}$ coincide with the standard one. By the uniqueness of tubular neighborhoods, we can make $N(D_{v})$ and $\theta^{v}_{\tau}$ standard as well.
@@
This sweep-out $\{\Sigma_{ts}\}_{(t,s)\in[0,2]\times B}$ satisfies the desired property. Indeed, if $\ell\subset B$ is a radial segment in $B$, $\{\Sigma_{2s}\}_{s\in\ell}$ appears in some $\{\Sigma^{k}_{s}\}_{s\in B}$ as a subfamily. Thus (2) and (3) hold. (1) is obvious from the construction. ∎
@@
After perturbing $Q$, we assume that $\partial Q$ intersects $\Sigma_{y}$ transversely. We can find a circle in $\partial Q\cap\Sigma_{y}$ that is essential in $\Sigma_{y}$. Indeed, if all the circles in $\partial Q\cap\Sigma_{y}$ were inessential, by the innermost disk argument, $\Sigma_{y}$ could be isotoped so that afterward $\Sigma_{y}\subset Q$. This is impossible because $\Sigma_{y}$ is a Heegaard surface and $M$ is not a $3$-sphere. So one of the circles in $\partial Q\cap\Sigma_{y}$ bounds a compressing disk for $\Sigma_{y}$. Define $D_{y}$ as such a disk. By definition, $D_{y}\cap D_{y_{j}}=\emptyset$ for $21\leq j\leq 2^{i+1}-$ 2 and $\mathrm{diam}\,\partial D_{y}<3^{1+\cdots+(i+1)}\cdot\delta$, which proves the claim. ∎
@@
If $d=2$, we can deduce a contradiction as follows. Let $V$ and $W$ be the handlebodies in $M$ bounded by $\Sigma$. Denote by $\Gamma_{V}(\Sigma)$ (resp. $\Gamma_{W}(\Sigma)$) the subcomplex of $\Gamma(\Sigma)$ spanned by compressing disks for $\Sigma$ that lie in $V$ (resp. $W$). Furthermore, denote by $\Gamma_{VW}(\Sigma)$ the union of all simplices that contain vertices in both $\Gamma_{V}(\Sigma)$ and $\Gamma_{W}(\Sigma)$. Thus, $\Gamma(\Sigma)=\Gamma_{V}(\Sigma)\cup\Gamma_{VW}(\Sigma)\cup\Gamma_{W}(\Sigma)$. Recall that the choice of $\varphi$ is arbitrary as long as it is homotopically nontrivial. By Claim 2.7 in [3], we can assume that $\varphi$ is represented by a loop $\gamma$ in $\Gamma(\Sigma)$ with the following properties: